# Constraining cosmological parameters using density split lensing and the conditional stellar mass function

### v1

@@ I Introduction
To infer cosmological parameters of the standard model of cosmology, called the $\Lambda$ Cold Dark Matter model ($\Lambda$CDM), previous works made use of the density-split statistics (DSS), which extracts information based on environments measured from smoothing foreground galaxy samples [1, 2, 3, 4]. Afterwards, these environments are assigned to different quantiles, which in turn are used to correlate with either foreground galaxies’ positions or background galaxies’ shear. This work focuses on the correlation with shear, which we call density split lensing (DSL). In [1, 2], cosmological and galaxy-halo parameters were inferred using the DSL, and they have shown that the DSL is a competitive method compared to the usually used second-order statistic. The reason is that second-order statistics can only extract the Gaussian information from the data. In contrast, DSS is a higher-order statistic and can extract information beyond the Gaussian part of the data. The potential power of higher-order statistics was nicely demonstrated in [5] for weak lensing statistics and in [6] for galaxy clustering statistics.
Although analytical models of the DSL [7, 8] provide insights into the underlying properties of the large-scale structure (LSS) they rely on several assumptions, like the galaxy bias, that are only valid if large scales are considered. In contrast, simulation-based models have the drawback that they are measured from expensive simulations and, therefore, depend on the noise of the simulations and the number of training nodes used to train the emulator. However, simulation-based models enable a more complicated connection between dark matter haloes and galaxies. A possible approach to model the galaxy-halo connection is using a Halo Occupation Distribution (HOD) [9], which populates the dark matter haloes with central and satellite galaxies depending on the halo’s mass. Using a HOD instead of linear galaxy bias, which is by construction only valid on large scales, allows the use of smaller scales.
@@
Similar to [4], we follow the idea of [18, 3, 4] and build a model based on the AbacusSummit [^1] https://abacussummit.readthedocs.io simulations [19]. Using the halo catalogues of the AbacusSummit simulations, we built a CSMF HOD that predicts the number of central and satellites per halo and the stellar mass of those galaxies [^2] A current implementation of the CSMF HOD will be made public here https://github.com/abacusorg/abacusutils.. Using those galaxies, we define environments that depend on the galaxy number density and measure those environments’ excess surface mass density (ESD). Based on different cosmologies and CSMF parameters, we built an emulator for the ESD, which can predict the model everywhere in the predefined parameters space. Emulators are a machine learning tool widely used in cosmology due to their speed and accuracy in interpolating the model vector between predefined training nodes [20, 21, 22, 23]. Furthermore, while we used the projected environment in [4], we follow here the work of [3] and define the environment in the three-dimensional space, which is only possible because we tailor our foreground galaxies towards the spectroscopic data from early data release of the Dark Energy Spectroscopic Instrument (DESI-EDR). In future work, we expect to measure the ESD from the Ultraviolet Near-Infrared Optical Northern Survey (UNIONS) [24], so here we take this as a reference survey for the background sources.
A natural extension to the work presented here would be to perform a similar analysis as in [25] and combine DSL with density split clustering as performed in [3]. As shown in [25], combining clustering and lensing statistics improves constraining cosmological and HOD parameter constraints. We ignored it for this work because we have seen in [25] that clustering statistics are significantly more challenging to emulate due to the small uncertainties.
This work is structured as follows. In Sect. II, we review the basics of the excess surface mass density (ESD). In Sect. III, we briefly describe the reference data used to define the prior range and create a realistic covariance matrix for a parameter forecast and validate the emulator’s accuracy. In Sect. IV, we review the basics of the AbacusSummit simulations that we use to build our model and the implementation of the CSMF. In Sect. VII, we describe our numerical covariance matrix estimation. In Sect. VIII, we describe the accuracy of our emulator model, and in Sect. IX, we perform a forecast and finalize our work in Sect. X.
@@ II Excess surface mass density
This work focuses on measuring the ESD of a geometrical thin lens, which is defined as
@@
where $\sum_{ls}$ is the sum over all lens source pairs. Since the ESD is insensitive to the source distance distribution, we used a single redshift slice to measure the ESD in the simulations. The measured ESDs with that approach give unbiased results when compared to real data ESD measurements, as long as the redshift distribution of the real data is behind the redshift distribution of the lens [26]. Therefore, this allows us to apply the derived model in this work to real data measurements without any modifications in the future. Although the so-called boost correction could potentially correct for an overlapping lens and source redshift distribution, a well-separated lens and source redshift distribution also avoid the complicated modelling of intrinsic alignments and magnification effects.
@@ III Reference data
UNIONS is a weak lensing survey that measures galaxy shapes and photometric redshifts above a declination of $+30^{\circ}$ and a galactic latitude $|b|>25^{\circ}$. In this work, the information we need to derive the model is the mean redshift $z\approx 0.65$ of the galaxies, which will be used to calibrate photometric redshift distributions. Besides this, we also use the shape measurements from the preliminary UNIONS catalogue (v1.3 ShapePipe) to estimate a covariance matrix.
DESI is a spectroscopic survey that seeks to map the LSS of the Universe by measuring the positions and redshifts of 30 million pre-selected galaxies across one-third of the night sky [29]. In this work, we make use of the BGS sample [11, 12] of the sv3 early-data release of DESI [30], where sv3 is a $140\,\mathrm{deg}^{2}$ of DESI using the same targeting as the main survey with a high fibre assignment completeness to enable unbiased small scale clustering measurements. In particular, we are using the clustering catalogues BGS_BRIGHT_N_clustering and BGS_BRIGHT_S_clustering, described in [31]. In addition, we extract the stellar masses and absolute magnitudes from the Legacy Survey DR9 in the $u,g,r$-bands [32, 33] from the fastspec-fuji-sv3-bright catalogue. These stellar masses are determined using the fastspecfit software described in [34]. Using the stellar masses and the $u-g$ colour of the galaxies, we select only red galaxies by applying the following criteria
@@
where the parameters are chosen such that we cut through the green valley of the galaxy sample, as seen in Fig. 1. Using this selection, we decreased the number of galaxies from 143853 to 62244. This sample of galaxies is used to estimate the stellar mass function (SMF), which in turn is used to define the prior range of our HOD parameters.
@@ IV AbacusSummit light cone simulations
This work uses the AbacusSummit [19] to build an emulator model to predict the ESD. Here, we mention only the essential details to follow the work of this paper. The AbacusSummit is a cosmological simulation run with the Abacus N-body code [35, 36]. The base AbacusSummit simulations follow the evolution of $6912^{3}$ dark matter particles in a periodic box with a side length of $2\,h^{-1}\text{Gpc}$, and a mass resolution of $2.1\times 10^{9}\,h^{-1}{\rm M_{\odot}}$. Group or haloes of particles are found with the CompaSO hybrid Friends-of-Friends/Spherical Overdensity algorithm [37]. Lightcone simulations are described in [37], which we use to construct realistic mock catalogues of the sky that vary with redshift. The light cones cover an octant of the sky up to a redshift of 0.8333At large redshift the octant decreases due to the placement of the boxes to build the light cones. and span 63 different cosmologies within an extended $w_{0}w_{a}$ parameter space (see Fig. 2),
@@
We use the shear information ($\kappa$ and $\gamma$) at a single redshift slice $z=0.65$ to measure the ESD described above. The chosen redshift is roughly the median redshift of the calibration sample used to calibrate the photometric redshift of the UNIONS data. The resolution of this shear map is $A_{\mathrm{pix}}=0.046\,\mathrm{arcmin}^{2}$ $(\texttt{nside}=16384)$, or a side length of $49.4\,h^{-1}\mathrm{kpc}$ at a redshift of $z=0.4$ given by the T17 cosmology.
@@
$M_{\mathrm{c}}^{*}({M_{\mathrm{h}}})=M_{0}\frac{({M_{\mathrm{h}}}/M_{1})^{\gamma_{1}}}{(1-{M_{\mathrm{h}}}/M_{1})^{\gamma_{1}-\gamma_{2}}}\ ,$ (IV.6)
@@ V Prior of the HOD parameters
The context regarding criterion 3 is that we save, due to storage issues, only $0.9$% of the particles available. To ensure that the satellite profile and the assembly bias modulation are possible, we ensured that none of the parameters requires more than $\sim 0.3\%$ of the available particles.
@@
where $a=2.0404$, $b=1.8140$ and $c=11.4499$. In Fig. 4, we show the distribution of stellar mass versus the spectroscopic redshift of the DESI-EDR and the stellar mass completeness function $M^{\mathrm{lim}}_{\star}$. To measure the SMF for $1210.2<\log_{10}[M_{\star}/(h^{-1}\mathrm{M}_{\odot})]<$ 12 with $\Delta\log_{10}[M_{\star}/(h^{-1}\mathrm{M}_{\odot})]=0.1$ and $0.40.1<z<$ 0.4 from those galaxies that fulfill the stellar mass completeness criteria, we follow [17] and compute for each galaxy its observable volume, where we used the fiducial cosmology c000 of AbacusSummit to measure comoving distances. The SMF is given by binning the stellar masses and summing over all these volumes, weighted by the weight per galaxy provided in the DESI-EDR catalogues for completeness to fibre collisions. Furthermore, we ignore that the DESI samples are missing objects that, although they appear brighter than $19.5$, have an apparent fibre magnitude too faint to properly measure redshifts. This reduces the completeness to $98\%$ [12]. However, even if we multiply the derived SMF with $1/0.8$, which is according to figure 12 of [12], the worst case, the measured SMF is still inside the allowed prior range (see black dots in Fig. IV.10). The SMF uncertainty is estimated by dividing the sample into ten subsamples. To predict SMF, we use Eq. IV.10, where the halo mass function $n({M_{\mathrm{h}}})$ is measured from the 25 phases of the AbacusSummit and the assumed cosmology is the fiducial cosmology c000 of AbacusSummit. We fixed the cosmology and then allowed only those HOD parameters with a $\chi^{2}<100$, where we used only the diagonal of the covariance matrix.
This threshold of 100 is arbitrarily chosen; however, as shown in Fig. 5, the resulting SMF predicted with the more sophisticated approach described in Sec. VI spread significantly more than the DESI-EDR measurement suggests. Again, we note that the estimate of the SMF uncertainty, the threshold of 100, and the simplified prediction with a non-varying cosmology are solely used to define the prior range of the HOD parameter space and are not used for the parameter forecasts. We have checked that using a different cosmology has not changed the prior range significantly.
@@
Value $10.6441$ $10.7807$ $0.2509$ $26.7791$ $0.1900$ $11.5309$ $1.7706$ $-0.5379$ $0.0$ $0.0$ $0.0$
To ensure that the parameters cover the whole valid prior range, we ran an MCMC process with 1000 walkers, where the log-likelihood is always -0.5 if the criteria are fulfilled. If the walkers have been running for long enough, they should cover the whole prior range. From those points in the chain, we randomly selected 6300 (100 per cosmology) combinations as points in parameter space, which we use to train our emulator. The resulting parameter space is shown in Fig. 6, where it is seen how correlated some of the parameters are. The assembly bias and satellite modulation parameter have fixed priors motivated from [4]. The lower bounds of $b_{0}$, $b_{1}$, and $\gamma_{2}$ and the upper bounds of $\gamma_{1}$ are similar to the ones used in [17].
@@ VI Modelling the summary statistics
Figure 7: Illustration of all ESD data vectors. The left panel shows the ESD around random query points of the highest (D10) and lowest (D1) decile. The middle and right panel shows the ESD around all lens galaxies in red or for the three distinct deciles. The shade region is the expected error estimated from the T17, and the error bars are estimated from the 25-phase realization given in AbacusSummit.
@@
To model the SMF for $1210.5<\log_{10}(M_{\star}/(h^{-1}\mathrm{M}_{\odot}))<$ 12777The lower bound is 10.5 instead of 10.2, which was used for the SMF measurements to determine the prior range, since we initially aimed to have volume-limited samples for a better comparison with the literature. However, we dropped that criterion and used all galaxies above $M_{\star}^{\mathrm{lim}}$. As long as we treat real data and simulations the same way, this is a fair approach. with $\Delta\log_{10}(M_{\star}/(h^{-1}\mathrm{M}_{\odot}))=0.1$ and $0.40.1<z<$ 0.4, we use again Eq. IV.10, where the halo mass function $n({M_{\mathrm{h}}})$ is modelled using the function form of [47] implemented in pyccl, which changes with cosmology and redshift. To include the redshift dependence, we measured the median redshift of the T17 mocks described below over all realizations for each stellar mass bin. Furthermore, to account for the cosmology dependence of potential SMF measurement, we need to modify the SMF prediction $\Phi^{\mathrm{model}}_{\mathrm{X}}(M_{\star})$ with
@@
In addition to the raw ESD around all the lens galaxies, we aim to measure how the ESD depends on the local galaxy density environments. Motivated by the previous works of [3], we define the three-dimensional environments by smoothing the three-dimensional galaxy distribution, $n_{\mathrm{g}}$, with a Gaussian filter of $R_{G}=8\ h^{-1}\mathrm{Mpc}$. We note that we used all galaxies with ${\log_{10}[M_{\star}/(h^{-1}\mathrm{M}_{\odot})]>10.5}$ and $0.30.1<z<$ 0.3 that fulfill the completeness criteria to define $n_{\mathrm{g}}$. The redshift cut of $z<0.3$ results from the fact that above that limit, the number density drops below $\sim 10$ galaxies per sphere of radius $2\times 8\ h^{-1}\mathrm{Mpc}$, and therefore, would result in noisy estimates of the environment. To ensure consistency between the ESD measurements, we also performed the same redshift cut for the ESD measurements of all galaxies. To account for masking, we divide the smoothed galaxy distribution, $n_{\mathrm{g}}^{R_{G}}$, with a smooth version of random points, $n_{\mathrm{r}}^{R_{G}}$, which are distributed in the same volume as the galaxies but several times denser. This corrected density environment,

### v2

@@ I Introduction
[…]
Although analytical models of the DSL [7, 8] provide insights into the underlying properties of the large-scale structure (LSS) they rely on several assumptions, like the galaxy bias, that are only valid if large scales are considered. In contrast, simulation-based models have the drawback that they are measured from expensive simulations and, therefore, depend on the noise of the simulations and the number of training nodes used to train the emulator. However, simulation-based models enable a more complicated connection between dark matter haloes and galaxies. A possible approach to model the galaxy-halo connection is using a Halo Occupation Distribution (HOD) [9], which populates the dark matter haloes with central and satellite galaxies depending on the halo’s mass. Using a HOD instead of linear galaxy bias, which is only valid on large scales, allows using smaller scales.
@@
Similar to [4], we follow the idea of [18, 3, 4] and build a model based on the AbacusSummit [^1] https://abacussummit.readthedocs.io simulations [19]. Using the halo catalogues of the AbacusSummit simulations, we built a CSMF HOD that predicts the number of central and satellites per halo and the stellar mass of those galaxies [^2] A current implementation of the CSMF HOD will be made public here https://github.com/abacusorg/abacusutils.. Using those galaxies, we define environments that depend on the galaxy number density and measure those environments’ excess surface mass density (ESD). Based on different cosmologies and CSMF parameters, we built an emulator for the ESD, which can predict the model everywhere in the predefined parameters space. Emulators are a machine learning tool widely used in cosmology due to their speed and accuracy in interpolating the model vector between predefined training nodes [20, 21, 22, 23, 24]. Furthermore, while we used the projected environment in [4], we follow here the work of [3] and define the environment in the three-dimensional space, which is only possible because we tailor our foreground galaxies towards the spectroscopic data from early data release of the Dark Energy Spectroscopic Instrument (DESI-EDR). In future work, we expect to measure the ESD from the Ultraviolet Near-Infrared Optical Northern Survey (UNIONS) [25], so here we take this as a reference survey for the background sources.
A natural extension to the work presented here would be to perform a similar analysis as in [26] and combine DSL with density split clustering as performed in [3]. As shown in [26], combining clustering and lensing statistics improves constraining cosmological and HOD parameter constraints. We ignored it for this work because we have seen in [26] that clustering statistics are significantly more challenging to emulate due to the small uncertainties. We therefore use only the galaxy-galaxy lensing signal for lens galaxies in different environments combined with predictions of the stellar mass function (SMF).
This work is structured as follows. In Sect. II, we review the basics of the excess surface mass density (ESD). In Sect. III, we briefly describe the reference data used to define the prior range and create a realistic covariance matrix for a parameter forecast and validate the emulator’s accuracy. In Sect. IV, we review the basics of the AbacusSummit simulations that we use to build our model and the implementation of the CSMF. In Sect. VII, we describe our numerical covariance matrix estimation using the simulations described in [27, thereafter T17]. In Sect. VIII, we describe the accuracy of our emulator model, and in Sect. IX, we perform a forecast and finalize our work in Sect. X.
@@ II Excess surface mass density
This work focuses on measuring the ESD of a geometrical thin lens, which is defined assuming circular symmetry as
@@
where $\sum_{ls}$ is the sum over all lens source pairs of a bin with center $r_{\mathrm{p}}$. Since the ESD dependency on the source distance distribution is cancelled out by the $\Sigma_{\mathrm{crit}}$ division, we used a single redshift slice to measure the ESD in the simulations. The measured ESDs with that approach give unbiased results when compared to real data ESD measurements, as long as the redshift distribution of the real data is behind the redshift distribution of the lens [28]. Therefore, this allows us to apply the derived model in this work to real data measurements without any modifications in the future. Although the so-called boost correction could potentially correct for an overlapping lens and source redshift distribution, a well-separated lens and source redshift distribution also avoid the complicated modelling of intrinsic alignments and magnification effects.
@@ III Reference data
UNIONS is a weak lensing survey that measures galaxy shapes and photometric redshifts above a declination of $30^{\circ}$ and a galactic latitude $|b|>25^{\circ}$. In this work, the information we need to derive the model is the mean redshift $z\approx 0.65$ of the galaxies, which will be used to calibrate photometric redshift distributions. Besides this, we also use the shape measurements from the preliminary UNIONS catalogue (v1.3 ShapePipe) to estimate a covariance matrix, which we further discuss in Sect. VII.
DESI is a spectroscopic survey that seeks to map the LSS of the Universe by measuring the positions and redshifts of 30 million pre-selected galaxies across one-third of the night sky [31]. In this work, we make use of the BGS sample [11, 12] of the sv3 early-data release of DESI [32], where sv3 is a $140\,\mathrm{deg}^{2}$ of DESI using the same targeting as the main survey with a high fibre assignment completeness to enable unbiased small scale clustering measurements. In particular, we are using the clustering catalogues BGS_BRIGHT_N_clustering and BGS_BRIGHT_S_clustering, described in [33]. In addition, we extract the stellar masses and absolute magnitudes from the Legacy Survey DR9 in the $u,g,r$-bands [34, 35] from the fastspec-fuji-sv3-bright catalogue. These stellar masses are determined using the fastspecfit software described in [36]. Using the stellar masses and the ugu-g colour of the galaxies, we select only red galaxies by applying the following criteria
@@
where the parameters are chosen such that we cut through the green valley of the galaxy sample, as seen in Fig. 1. Using this selection, we decreased the number of galaxies from 143853 to 62244. This sample of galaxies is used to estimate the SMF, which in turn is used to define the prior range of our HOD parameters.
@@ IV AbacusSummit light cone simulations
This work uses the AbacusSummit [19] to build an emulator model to predict the ESD. Here, we mention only the essential details to follow the work of this paper. The AbacusSummit is a cosmological simulation run with the Abacus N-body code [37, 38]. The base AbacusSummit simulations follow the evolution of $6912^{3}$ dark matter particles in a periodic box with a side length of $2\,h^{-1}\text{Gpc}$, and a mass resolution of $2.1\times 10^{9}\,h^{-1}{\rm M_{\odot}}$. Group or haloes of particles are found with the CompaSO hybrid Friends-of-Friends/Spherical Overdensity algorithm [39]. Lightcone simulations are described in [39], which we use to construct realistic mock catalogues of the sky that vary with redshift. The light cones cover an octant of the sky up to a redshift of 0.8333At large redshift the octant decreases due to the placement of the boxes to build the light cones. and span 63 different cosmologies within an extended $w_{0}w_{a}$CDM parameter space (see Fig. 2),
@@
We use the reduced shear information, $g=\gamma/(1-\kappa)$, a single redshift slice $z=0.65$ to measure the ESD described above. The chosen redshift is roughly the median redshift of the calibration sample used to calibrate the photometric redshift of the UNIONS data. The resolution of this shear map is $A_{\mathrm{pix}}=0.046\,\mathrm{arcmin}^{2}$ $(\texttt{nside}=16384)$, or a side length of $49.4\,h^{-1}\mathrm{kpc}$ at a redshift of $z=0.4$ given by the T17 cosmology.
@@
$M_{\mathrm{c}}^{*}({M_{\mathrm{h}}})=M_{0}\frac{({M_{\mathrm{h}}}/M_{1})^{\gamma_{1}}}{(1+{M_{\mathrm{h}}}/M_{1})^{\gamma_{1}-\gamma_{2}}}\ ,$ (IV.6)
@@ V Prior of the HOD parameters
The context regarding criterion 3 is that we save, due to storage issues, only 0.90.9% of the particles available. To ensure that the satellite profile and the assembly bias modulation are possible, we ensured that none of the parameters requires more than $\sim 0.3\%$ of the available particles.
@@
where $a=2.0404$, $b=1.8140$ and $c=11.4499$. In Fig. 4, we show the distribution of stellar mass versus the spectroscopic redshift of the DESI-EDR and the stellar mass completeness function $M^{\mathrm{lim}}_{\star}$. To measure the SMF for $10.2<\log_{10}[M_{\star}/(h^{-1}\mathrm{M}_{\odot})]<12$ with $\Delta\log_{10}[M_{\star}/(h^{-1}\mathrm{M}_{\odot})]=0.1$ and $0.1<z<0.4$ from those galaxies that fulfill the stellar mass completeness criteria, we follow [17] and compute for each galaxy its observable volume, where we used the fiducial cosmology c000 of AbacusSummit to measure comoving distances. The SMF is given by binning the stellar masses and summing over all these volumes, weighted by the weight per galaxy provided in the DESI-EDR catalogues for completeness to fibre collisions. Furthermore, we ignore that the DESI samples are missing objects that, although they appear brighter than $19.5$, have an apparent fibre magnitude too faint to properly measure redshifts. This reduces the completeness to $98\%$ [12]. However, even if we multiply the derived SMF with $1/0.8$, which is according to figure 12 of [12], the worst case, the measured SMF is still inside the allowed prior range (see black dots in Fig. IV.10). The SMF uncertainty is estimated by dividing the sample into ten subsamples. To predict SMF, we use Eq. IV.10, where the halo mass function $n({M_{\mathrm{h}}})$ is measured from the 25 phase realizations of the AbacusSummit to reduce the effect of cosmic variance and the assumed cosmology is the fiducial cosmology c000 of AbacusSummit. We fixed the cosmology and then allowed only those HOD parameters with a $\chi^{2}<100$, where we used only the diagonal of the covariance matrix.
This threshold of 100 is arbitrarily chosen; however, as shown in Fig. 5, the resulting SMF predicted with the more sophisticated approach described in Sect. VI spread significantly more than the DESI-EDR measurement suggests. Again, we note that the estimate of the SMF uncertainty, the threshold of 100, and the simplified prediction with a non-varying cosmology are solely used to define the prior range of the HOD parameter space and are not used for the parameter forecasts. We have checked that using a different cosmology has not changed the prior range significantly.
@@
Value $10.6441$ $10.7807$ $0.2509$ $26.7791$ $0.1900$ $11.5309$ $1.7706$ 0.5379-0.5379 $0.0$ $0.0$ $0.0$
To ensure that the parameters cover the whole valid prior range, we ran an MCMC process with 1000 walkers, where the log-likelihood is always -0.5 if the criteria are fulfilled. If the walkers have been running for long enough, they should cover the whole prior range. From those points in the chain, we randomly selected 6300 (100 per cosmology) combinations as points in parameter space, which we use to train our emulator. The resulting parameter space is shown in Fig. 6, where it is seen how correlated some of the parameters are. The assembly bias and satellite modulation parameter have fixed priors motivated from [4]. The lower bounds of $b_{0}$, $b_{1}$, and $\gamma_{2}$ and the upper bounds of $\gamma_{1}$ are similar to the ones used in [17]. Although these prior restrictions are dominant for some parameters, as we will see in Sect. IX, they do not affect essential HOD parameters like $M_{0,1}$ or $\sigma_{\mathrm{c}}$.
@@ VI Modelling the summary statistics
Figure 7: Illustration of all ESD data vectors. The left panel shows the ESD around random query points of the highest (D10) and lowest (D1) decile. The middle and right panel shows the ESD around all lens galaxies in red or for the three distinct deciles. The shade region is the expected error estimated from the T17 mimicking a possible DESIxUNIONS setup as described in Sect. VII, and the error bars are estimated from the 25-phase realization given in AbacusSummit.
@@
To model the SMF for $10.5<\log_{10}(M_{\star}/(h^{-1}\mathrm{M}_{\odot}))<12$ [^7] The lower bound is 10.5 instead of 10.2, which was used for the SMF measurements to determine the prior range, since we initially aimed to have volume-limited samples for a better comparison with the literature. However, we dropped that criterion and used all galaxies above $M_{\star}^{\mathrm{lim}}$. As long as we treat real data and simulations the same way, this is a fair approach. with $\Delta\log_{10}(M_{\star}/(h^{-1}\mathrm{M}_{\odot}))=0.1$ and $0.1<z<0.4$, we use again Eq. IV.10, where the halo mass function $n({M_{\mathrm{h}}})$ is modelled using the function form of [48] implemented in pyccl, which changes with cosmology and redshift. To include the redshift dependence, we measured the median redshift of the T17 mocks described below over all realizations for each stellar mass bin. Furthermore, to account for the cosmology dependence of potential SMF measurement, we need to modify the SMF prediction $\Phi^{\mathrm{model}}_{\mathrm{X}}(M_{\star})$ with
@@
In addition to the raw ESD around all the lens galaxies, we aim to measure how the ESD depends on the local galaxy density environments. Motivated by the previous works of [3], we define the three-dimensional environments by smoothing the three-dimensional galaxy distribution, $n_{\mathrm{g}}$, with a Gaussian filter of $R_{G}=8\ h^{-1}\mathrm{Mpc}$. We note that we used all galaxies with $\log_{10}[M_{\star}/(h^{-1}\mathrm{M}_{\odot})]>10.5$} and $0.1<z<0.3$ that fulfill the completeness criteria to define $n_{\mathrm{g}}$. The redshift cut of $z<0.3$ results from the fact that above that limit, the number density drops below $\sim 10$ galaxies per sphere of radius $2\times 8\ h^{-1}\mathrm{Mpc}$, and therefore, would result in noisy estimates of the environment. To ensure consistency between the ESD measurements, we also performed the same redshift cut for the ESD measurements of all galaxies. To account for masking, we divide the smoothed galaxy distribution, $n_{\mathrm{g}}^{R_{G}}$, with a smooth version of random points, $n_{\mathrm{r}}^{R_{G}}$, which are distributed in the same volume as the galaxies but several times denser. This corrected density environment,